McDormand, Inc., reported a $3,500 unfavorable price variance for variable overhead and a $35,000 unfavorable price variance for fixed overhead. The flexible budget had $1,084,500 variable overhead based on 36,150 direct labor-hours; only 34,110 hours were worked. Total actual overhead was $1,829,800. The number of estimated hours for computing the fixed overhead application rate totaled 38,400 hours.  
  

Required:

a. Prepare a variable overhead analysis.
b. Prepare a fixed overhead analysis.
Answer #1

RA) price variance= 3,500 U  answer

or 1,026,800- 1,023,300= 3,500 U


1,084,500/ 36,150= 30      34,110* 30= 1,023,300


efficiency variance= 1,084,500- 1,023,300= 61,200 F answer

variable OH variance= 61,200- 3,500= 57,700 F answer


RB) 

price variable= 35,000 U answer


step 1) 1,829,800- 1,026,800= 803,000

803,000- 35,000= 768,000

768,000/ 38400= 20

20* 36,150= 723,000

price variance  803,000- 768,000= 35,000 U

production volume variance  768,000- 723,000= 45,000 U answer

fixed overhead variance  35,000+ 45,000= 80,000 U answer
Answer #2

1,084,500- 1,026,800= 57,700 


1,084,500- 57,700 (variable OH variance)= 1,026,800